# The SOFIA Massive (SOMA) Radio Survey. II. Radio Emission from High-Luminosity Protostars

### v1

@@ 3 Results
a Inner scale corresponds to source 30 that we are resolving into sources 30N and 30S (see section 3.1.3 and Table 14 for more details).
Scale | R.A | Decl. | Spectral | ||||
(J2000) | (J2000) | (mJy) | (mJy) | (mJy) | (mJy) | Index | |
SOMA | 16:59:41.63 | 40.03.43.60 | 18.92(1.9) | 23.50(2.4) | 34.27(3.7) | 31.56(3.5) | 0.4(0.1) |
Intermediate | 16:59:42.17 | 40.03.46.85 | 19.66(2.0) | 27.63(2.8) | 31.84(3.2) | 33.34(3.4) | 0.4(0.1) |
Inner | 16:59:41.63 | 40.03.43.73 | 8.18(0.87) | 13.40(1.43) | 20.80(2.19) | 22.90(2.43) | 0.8(0.1) |
Scale | R.A | Decl. | Spectral | ||||
(J2000) | (J2000) | (mJy) | (mJy) | (mJy) | (mJy) | Index | |
SOMA | 16:59:43.01 | 40.03.11.56 | 0.39 | 0.47(0.22) | 0.59(0.64) | 0.87(0.66) | 0.7 |
Intermediate | |||||||
Inner | 0.06 | 0.06 | 0.09 | 0.08 | 0.4 |
Scale | R.A | Decl. | Spectral | ||||
(J2000) | (J2000) | (mJy) | (mJy) | (mJy) | (mJy) | Index | |
SOMA | 13:11:10.49 | 62.34.38.80 | 16.57(2.11) | 12.98(2.36) | 3.42(2.63) | 29.72 | -1.0 |
Intermediate | 13:11:10.52 | 62.34.39.08 | 7.23(0.83) | 4.55(0.77) | 1.59(0.83) | 9.33 | -1.2 |
Inner | 13:11:10.56 | 62.34.38.31 | 3.39(0.13) | 1.61(0.89) | 1.37(0.19) | 1.92(0.38) | -0.7(0.2) |
Scale | R.A | Decl. | Spectral | ||||
(J2000) | (J2000) | (mJy) | (mJy) | (mJy) | (mJy) | Index | |
SOMA | 13:11:13.64 | 62.34.40.70 | 1.25(2.37) | 1.05(2.14) | 0.71(2.18) | 3.22(8.43) | -0.3(3.1) |
Intermediate | |||||||
Inner | 13:11:13.75 | 62.34.41.31 | 0.96(0.01) | 0.74(0.04) | 0.91(0.09) | 2.70(0.27) | 0.2(0.2) |
Scale | R.A | Decl. | Spectral | ||||
(J2000) | (J2000) | (mJy) | (mJy) | (mJy) | (mJy) | Index | |
SOMA | 19:23:06.61 | +14.20.12.00 | 44.30(4.54) | 52.93(5.44) | 26.11(3.34) | 40.86(4.50) | -0.3(0.1) |
Intermediate | 19:23:06.69 | +14.20.11.22 | 0.21(0.04) | 0.01(0.04) | 0.33(0.08) | 0.34(0.07) | 0.7(0.5) |
Inner | 19:23:06.66 | +14.20.11.66 | 0.04(0.01) | 0.03(0.01) | 0.27(0.02) | 0.35(0.02) | 1.6(0.2) |
Scale | R.A | Decl. | Spectral | ||||
(J2000) | (J2000) | (mJy) | (mJy) | (mJy) | (mJy) | Index | |
SOMA | 16:52:04.67 | 46.08.34.20 | 5.50(0.71) | 3.64(0.68) | 4.11(2.16) | 4.27(4.61) | -0.6(0.5) |
Intermediate | 16:52:04.62 | 46.08.36.04 | 5.31(0.57) | 6.86(0.73) | 4.58(1.04) | 4.03(2.05) | -0.1(0.2) |
Inner | 16:52:04.94 | 46.08.31.32 | 0.86(0.09) | 1.51(0.16) | 2.44(0.26) | 3.00(0.33) | 0.9(0.1) |
@@
Note. — The reported values correspond to a circle of radius R for the SOMA scale and a box of height h and width w for the Intermediate scale. In some cases the inner scale between the C and K images are different so we report the larger area that we consider any source of emission within the inner region, we also have two different types of regions:
@@
b Scale corresponds to an ellipse of major axis a and minor axis.
@@ 3.1 Morphology and Multiplicity
All the target regions presented in this paper have been observed in the centimeter continuum. We describe source morphology as “compact” if the detection shows no structure on the scale of a few synthesized beams or greater. Otherwise we describe it as “extended”. Below, we describe the centimeter wavelength detections toward each target; for a detailed literature review on each of these regions, see SOMA II.
@@
G45.12+0.13 is a massive star-forming region, also know as IRAS 19111+1048, that is associated with an UC HII region. It is located at 7.4 kpc (Ginsburg et al., 2011) and is part of the Galactic Ring Survey Molecular Cloud (GRSMC) G45.46+0.05, a large star formation complex (Wood & Churchwell, 1989; Simon et al., 2001).
G45.12+0.13 has been analyzed at multiple wavelengths ranging from IR, sub-mm and radio (Hunter et al., 1997; Vig et al., 2006; Azatyan et al., 2022). Hunter et al. (1997) reported bipolar outflows towards this region using CO () maps. From this study, the highest velocity outflow is centered around a source termed S14, while an additional outflow is centered around another UC HII region source G45.12+0.13 West. Both of these UC HII regions contain type-I OH masers (Argon et al., 2000). Additionally, with radio observations from the Giant Metrewave Radio Telescope (GMRT) at 1280, 610 and 325 MHz, Vig et al. (2006) suggested that this region shows a highly homogeneous ionized medium morphology and SOMA II confirm this morphology with their MIR images. Furthermore, Vig et al. (2006) concluded that G45.12+0.13 contains a cluster of ZAMS stars energizing a compact and evolved HII region. In total they reported 20 sources, with S14 being the central UC HII source and S20 a non-thermal source according to its radio emission.
From our VLA observations at 1.3 and 6 cm we report two detections corresponding to sources S14 and S20 from Vig et al. (2006). Both detections have an offset in the peak intensity position from the results of Vig et al. (2006) using GMRT observations, by a difference of around 2.1′′ to the SE (see Figure 2). This systematic offset may be due to astrometric uncertainties in the GMRT data.
It is important to highlight that despite our data being of higher resolution, we only detect two of the twenty compact sources from the cluster reported by Vig et al. (2006). From our data we infer that most of the emission to the northern side of the central UC HII region arises from an extended and diffuse irregular cloud of ionized material instead of a cluster of compact sources.
In order to confirm our results, we produce an image using shorter baseline observations from the VLA archive (project code 20A-519; PI: David Neufeld) at L-band (1.5 GHz) in order to compare the results with the 1.28 GHz image from Vig et al. (2006). Figure 2 shows this L-band continuum image with overlaid C, K and L-band VLA contours. The positions of the cluster of point sources reported by Vig et al. (2006) at 1280 MHz are marked with a white symbols.
@@
From our VLA observations, source S14 is located at R.A.(J2000) , Decl.(J2000) with a spectral index of 0.5. S14 presents a very complex morphology, and in fact at the shorter frequency (L-band) it appears to have an hour-glass morphology, as seen in the right plot of Figure 2. This observed morphology is similar to the one detected towards source G45.47+0.05 by Zhang et al. (2019a), who interpreted this as arising from a photoionized outflow. Furthermore, the molecular outflows observed towards G45.12+0.13 (see Figure 1) are consistent with the direction of the axis of the hour-glass of source S14. Meanwhile S20 is located at R.A.(J2000) , Decl.(J2000) with a spectral index of 0.9, a value that is consistent with the non-thermal detection made by Vig et al. (2006). S20 presents a very compact morphology at both frequency bands.
@@
G309.92+0.48 or IRAS 13471-6120 is a region located at 5.5 kpc (Murphy et al., 2010). It is associated with OH and methanol masers, as well as radio continuum emission from an UC HII region at 8.6 and 19 GHz (Caswell, 1997; Phillips et al., 1998; Walsh et al., 1998). In the MIR, G309.92+0.48 is resolved into several sources: De Buizer et al. (2000) reported it to be composed of a total of 6 point sources (labeled 1 to 6) using high resolution (′′) 11.7 m data from Gemini (see also SOMA II).
From our ATCA observations, we have detected two sources that we name ATCA1 and ATCA2 (see Figure 1). Source ATCA1 is detected at all frequencies (5.5, 9, 17, and 22.8 GHz) and source ATCA2 is only detected at the lower frequencies (5.5 and 9 GHz). Both sources are consistent with the previously reported sources by Phillips et al. (1998) at 8.6 GHz, sources that are associated with the MIR sources 1 and 2. Murphy et al. (2010) presented radio continuum observations at 20 GHz towards this region, detecting only source ATCA1, which is consistent with our results of non-detection for source ATCA2 at higher frequencies.
@@
For ATCA1, Murphy et al. (2010) presented a spectral index of 1.2 between 20 GHz and 0.843 GHz. Our estimated spectral index value is different (0.7), but still both results fall into the range of a typical UC HII region. ATCA2 shows non-thermal emission given the negative value in the spectral index. No molecular outflows have been detected towards this region.
@@
G35.58-0.03 (hereafter G35.58) is a star-forming region located at a kinematic distance of 10.2 kpc (Watson et al., 2003). G35.58 has been resolved into two UC HII regions (from now on called sources 30 and 31) using 2 and 3.6 cm observations, where water and OH masers have been detected towards source 30 (Kurtz et al., 1994; Caswell et al., 1995). Additionally, Zhang et al. (2014a) shows evidence of an ionized outflow driving a NE-SW molecular outflow centered in source 30.
@@
30A | 18:56:22.52 | +02.20.30.5 | 0.7(0.1) | IR | Yesa |
30B | 18:56:22.62 | +02.20.34.7 | -1.1 | Yesb | |
30C | 18:56:22.57 | +02.20.24.3 | -0.3 | Yesb | |
30D | 18:56:22.47 | +02.20.24.8 | -1.2 | Yesb | |
30E | 18:56:22.55 | +02.20.21.4 | -0.5(0.1) | Yesb | |
30F | 18:56:22.43 | +02.20.20.5 | -0.5 | Yesb | |
30G | 18:56:22.39 | +02.20.19.7 | -1.0 | Yesb | |
30H | 18:56:22.46 | +02.20.18.9 | -1.9 | Yesb | |
30J | 18:56:22.19 | +02.20.21.5 | 0.6 | Yesb | |
30K | 18:56:22.19 | +02.20.21.5 | 0.1 | Yesb | |
30N | 18:56:22.53 | +02.20.27.6 | -0.5(0.4) | IR, mm | Yesc |
30S | 18:56:22.52 | +02.20.27.2 | 0.2(0.1) | IR, mm | Yesc |
31 | 18:56:22.62 | +02.20.27.9 | 0.0(0.1) | MIR, mm | No |
@@
G35.58-0.03 30 and 31: Sources 30 and 31 have previously been referred to as UC HII regions (e.g., Kurtz et al., 1994). For source 31, based on its morphology in the centimeter and its estimated spectral index, we agree with its nature being an UC HII region. Moreover, source 31 appears to have already a central protostar based on its association with a bright and compact MIR source (SOMA II).
To our knowledge, we are resolving for the first time source 30 into two components. Specifically, at the C-band (see Figure 3) source 30 appears slightly resolved into two sources, which we have labeled 30S and 30N and they appear to be slightly aligned in the NE-SW direction.
Their elongation, which appears aligned in the direction of the larger scale molecular outflow, and their association with water masers appears to suggest a scenario in which source 30S and 30N could be part of an ionized jet. However, we cannot be certain of this with the current observations. Moreover, source 30, which shows several tracers of ongoing star formation, appears to be in a younger evolutionary stage than source 31. Furthermore, the estimated spectral index of 0.7, derived from the inner scale encompassing Source 30, aligns with the typical values observed for ionized jets associated with young stellar objects (YSOs). (e.g., Reynolds (1986); Anglada et al. (1998); Tanaka et al. (2016))
@@ 3.1.4 IRAS 16562-3959
IRAS 16562-3959, also know as G345.49+1.47, is a region located at 1.7 kpc and the massive core host of a massive star in an early stage of evolution (Guzmán et al., 2010). This region has been observed at multiple wavelengths, including centimeter continuum, mm, IR and X-ray (Guzmán et al., 2010, 2014, 2016, 2018; Liu et al., 2019; Montes et al., 2020).
@@
Guzmán et al. (2011) reported a quadrupolar molecular outflow using CO(6-5) and CO(7-6) lines. This morphology is explained by the presence of two collimated bipolar outflows, one lying in the SE–NW direction related to the ionized jet detected by Guzmán et al. (2010), and another in the N–S direction that could be related to the unresolved millimeter source 13 in Guzmán et al. (2014).
@@
Guzmán et al. (2016) reported nine radio detections at 5, 9, 17, and 19 GHz using ATCA observations. The central source has a counterpart at 3 mm and X-ray (source 10 at 3 mm and source 161 in X-ray) (Guzmán et al., 2014; Montes et al., 2020), and has also been detected in the MIR by the SOMA II study.
@@
For our analysis we used ATCA images at 5, 9, 17 and 19 as presented by Guzmán et al. (2016). We focus on regions IRAS 16562-3959 and 16562-3959 N. With IRAS 16562-3959 N being a new region added to the sample in SOMA III, due to the detection of a point source in the IR maps, by the observations analyzed in SOMA II.
The SOMA scale in IRAS 16562-3959 (see Figure 1) includes only the main source (C), the two inner lobes (i.e., I-E and I-W) and the three additional sources (1 to 3) reported by Guzmán et al. (2016). So, for the purpose of this study, we only focus on the sources within the SOMA scale. In Table 15 we present more information about the multiplicity in IRAS 16562-3959 (within the SOMA scale) such as the R.A. and Decl. positions, the estimated spectral indices, association with other wavelengths, and whether or not the source is a newly reported radio detection.
C | 16:59:41.61 | 40:03:43.7 | 0.8(0.1) | Xray, IR, mm | No |
I-E | 16:59:41.93 | 40:03:44.6 | 0.6(0.1) | Xray, mm | No |
I-W | 16:59:41.32 | 40:03:43.1 | 0.6(0.1) | Xray, mm | No |
1 | 16:59:42.42 | 40:03:31.0 | 1.5 | No | |
2 | 16:59:42.86 | 40:03:36.3 | 1.9 | Xray, IR, mm | No |
3 | 16:59:41.67 | 40:03:47.9 | 3.3 | No | |
4 | 16:59:40.77 | 40:03:46.9 | 1.3 | Yesa |
a These sources have been detected in Guzmán et al. (2010), but were not reported. For completeness we include them following our criteria of 5 detections.
IRAS 16562-3959: We find 7 radio sources within the SOMA scale. Detections C, I-E and I-W (following the nomenclature of Guzmán et al. (2010) and Guzmán et al. (2016)) were detected at all frequency bands. Meanwhile detections 1, 2 and 3 were only seen at the higher frequency bands (17.0 and 22.8 GHz) and a new detection, labeled as detection 4 (following the convention from Guzmán et al. (2016)), was found only at the lower frequency bands (5.5 and 9.0 GHz).
Detections C, 1-E and 1-W all show a compact morphology and using our own regions to measure the flux densities, we report an estimated spectral index of 0.8 for detection C. This value is consisted with the spectral index reported by Guzmán et al. (2014) of 0.920.01 at a wide frequency range of 1 to 100 GHz. Both our calculation and the previous reported values are in the range for typical ionized jets associated with YSOs (e.g., Reynolds (1986); Anglada et al. (1998); Tanaka et al. (2016)), and is consisted with the contemporaneous radio jet/HC HII region nature described in detail in Guzmán et al. (2016).
Detections 1-E and 1-W are considered lobes of the radio-jet in source C, and for both we report negative spectral indices which suggest non-thermal emission. The location of these two detections, the morphology and the optically thin free–free emission suggest that radio emission of the lobes comes from the synchrotron radiation that arises from the ionized radio jet. In contrast with Guzmán et al. (2016), we report some differences in the spectral index values due to the different regions used to measure the flux densities, but the non-thermal nature of the emission in the lobes remains.
@@
Given the clear jet-like morphology we also measure the fluxes using an intermediate scale, as previously described, that encompass the central source (detection C) and the two radio lobes. From this scale we report a spectral index of 0.4.
@@
Sources 1, 2 and 3 present a compact morphology and we report lower limits in spectral index shown in Table 15. We have similar results to those of Guzmán et al. (2016) so arrive at the same conclusions in terms of the nature of the sources. Guzmán et al. (2016) explains the nature of these detections in Section 4.2.2 of their paper. In summary, source 1 may be of extragalactic nature, source 2 is believed to be an HC HII region associated with a young high-mass star, and the free–free emission from source 3 comes from a low-mass YSO.
@@
From our criteria and for full completeness we report all detections above a 5 threshold. As a result of this, we report a new detection, source 4, that is only detected at C-band and shows a small elongation to the South (see Figure 1). We report an upper limit in the estimated spectral index shown in Table 15. This source can be clearly seen in Figure 7 of Guzmán et al. (2016), but they do not report this as a detection.
@@
From previous radio continuum observations toward G305.200.21 there are reports of two detections, which are prominent extended HII regions labeled as G305HII(SE) and G305HII by Walsh et al. (2007). These regions are located about 30′′ to the southeast of G305A and 15′′ to the southwest of G305B, respectively (Walsh et al., 2007; Phillips et al., 1998).
@@
G305A (G305.21+0.21): This source is prominent at MIR and FIR (SOMA II) and has been observed at 8.6 and 18 GHz, but no radio detections have been reported at sensitivity levels of 0.9 mJy and 0.22 mJy, respectively (Phillips et al., 1998; Walsh et al., 2007). With our ATCA observations at 5.5, 9, 17 and 22.8 GHz, we report a single detection within the region G305A that we label as G305A ATCA-1 (see Figure 1: G305.21+0.21 (G305A)). This source coincides with the methanol maser reported by Walsh & Burton (2006) and shows a slight elongation at the lower frequencies.
For G305A we report an estimated spectral index of 0.3, a result that indicates thermal emission. SOMA II suggested that G305A is a much younger and more embedded source than G305B and in a hot core phase, rich in molecular tracers (Walsh et al., 2007).
G305B (G305.20+0.21): Region G305B is the brightest MIR source associated with the methanol masers. However, no radio emission has been previously detected towards this region (Walsh et al., 2007). In our ATCA observations (see Figure 1) we report 4 sources labeled G305B ATCA-1, G305B ATCA-2, G305B ATCA-3 and G305B ATCA-4 (hereafter sources 1, 2, 3 and 4). In Table 16 we present more information about the multiplicity in G305B such as the R.A. and Decl. positions, the estimated spectral index, association with other wavelengths, and whether this source is a new radio detection or not.
@@
G305B ATCA-1 | 13:11:10.56 | 62.34.38.3 | -0.7(0.2) | IR | Yesa |
G305B ATCA-2 | 13:11:10.04 | 62.34.40.1 | -0.8(0.2) | Yesb | |
G305B ATCA-3 | 13:11:09.63 | 62.34.38.7 | -0.6(0.2) | Yesb | |
G305B ATCA-4 | 13:11:10.99 | 62.34.37.8 | -1.7 | Yesb |

### v3

@@ 3 Results
Inner scale corresponds to source 30 that we are resolving into sources 30N and 30S (see section 3.1.3 and Table 6 for more details).
The following are ATCA observations:
a The frequencies at which the flux was measured were 5.5, 9.0, 17.0 and 22.8 GHz.
b The frequencies at which the flux was measured were 5.0, 9.0, 17.0 and 19.0 GHz.
c The frequencies at which the flux was measured were 5.0, 7.0, 20.4 and 24.4 GHz.
@@
Note. — The reported values correspond to a circle of radius R for the SOMA scale and a box of height h and width w for the Intermediate scale. In some cases, the inner scales of the C- and K-band images differ. In these situations, we report the larger area and consider any emission within the inner region. We also distinguish between two types of regions:
@@
b Scale corresponds to an ellipse of major axis a and minor axis b.
@@ 3.1 Morphology and Multiplicity
All the target regions presented in this paper have been observed in the centimeter continuum. Our criteria for multiplicity is the presence of two or more spatially resolved 5 detections. Furthermore, we describe source morphology as “compact” if the detection shows no structure on the scale of a few number of synthesized beams or greater. Otherwise we describe it as “extended”. Below, we describe the centimeter wavelength detections toward each target; for a detailed literature review on each of these regions, see SOMA II.
@@
G45.12+0.13 is a massive star-forming region, also know as IRAS 19111+1048, that is associated with an UC HII region. It is located at 7.4 kpc (A. Ginsburg et al., 2011) and is part of the Galactic Ring Survey Molecular Cloud (GRSMC) G45.46+0.05, a large star formation complex (D. O. S. Wood & E. Churchwell, 1989; R. Simon et al., 2001).
G45.12+0.13 has been analyzed at multiple wavelengths ranging from IR, sub-mm and radio (T. R. Hunter et al., 1997; S. Vig et al., 2006; N. Azatyan et al., 2022; A. Y. Yang et al., 2021). T. R. Hunter et al. (1997) reported bipolar outflows towards this region using CO () maps. From this study, the highest velocity outflow is centered around a source termed S14, while an additional outflow is centered around another UC HII region source G45.12+0.13 West. Both of these UC HII regions contain type-I OH masers (A. L. Argon et al., 2000). Additionally, with radio observations from the Giant Metrewave Radio Telescope (GMRT) at 1280, 610 and 325 MHz, S. Vig et al. (2006) suggested that this region shows a highly homogeneous ionized medium morphology and SOMA II confirm this morphology with their MIR images. Furthermore, S. Vig et al. (2006) concluded that G45.12+0.13 contains a cluster of ZAMS stars energizing a compact and evolved HII region. In total they reported 20 sources, with S14 being the central UC HII source and S20 a non-thermal source according to its radio emission.
From our VLA observations at 1.3 and 6 cm we report two detections corresponding to sources S14 and S20 from S. Vig et al. (2006). Both detections have an offset in the peak intensity position from the results of S. Vig et al. (2006) using GMRT observations, by a difference of around 2.1′′ to the SE (see Figure 2). This systematic offset may be due to astrometric uncertainties in the GMRT data which we do not have enough information to quantify. Contributions to the uncertainty are expected to include calibrator position, phase transfer and self-calibration. The peak can also change with resolution given the presence of extended components.
It is important to highlight that despite our data being of higher resolution, we only detect two of the twenty compact sources from the cluster reported by S. Vig et al. (2006). From our data we infer that most of the emission to the northern side of the central UC HII region arises from an extended and diffuse irregular cloud of ionized material instead of a cluster of compact sources. The presence of a position offset does not change this conclusion.
In order to confirm our results, we produce an image using shorter baseline observations from the VLA archive (project code 20A-519; PI: David Neufeld) at L-band (1.5 GHz) in order to compare the results with the 1.28 GHz image from S. Vig et al. (2006). Figure 2 shows this L-band continuum image with overlaid C, K and L-band VLA contours. The positions of the cluster of point sources reported by S. Vig et al. (2006) at 1280 MHz are marked with a white symbols.
@@
From our VLA observations, source S14 is located at R.A.(J2000) , Decl.(J2000) with a spectral index of 0.5. S14 presents a very complex morphology, and in fact at the shorter frequency (L-band) it appears to have an hour-glass morphology, as seen in the right plot of Figure 2. This observed morphology is similar to the one detected towards source G45.47+0.05 by Y. Zhang et al. (2019b), who interpreted this as arising from a photoionized outflow. Furthermore, the molecular outflows observed towards G45.12+0.13 (see Figure 1) are consistent with the direction of the axis of the hour-glass of source S14.
Meanwhile S20 is located at R.A.(J2000) , Decl.(J2000) , presents a very compact morphology at both frequency bands and a spectral index of 0.9, a value that is consistent with the non-thermal detection made by S. Vig et al. (2006). Moreover, S. Vig et al. (2006) concluded that, even though source S20 is in the vicinity of G45.12+0.13, it is not likely to be associated with the UCHII region. If this source is indeed independent, the nature of this source could be extragalactic, but more information and observations to confirm this assessment are needed.
@@
G309.92+0.48 or IRAS 13471-6120 is a region located at 5.5 kpc (T. Murphy et al., 2010). It is associated with OH and methanol masers, as well as radio continuum emission from an UC HII region at several radio observations ranging from 5 to 24 GHz (J. L. Caswell, 1997; C. J. Phillips et al., 1998; A. J. Walsh et al., 1998; A. L. Patel et al., 2025). A. Y. Yang et al. (2021, 2022) reported a CO molecular outflow towards this region. In the MIR, G309.92+0.48 is resolved into several sources: J. M. De Buizer et al. (2000) reported it to be composed of a total of 6 point sources (labeled 1 to 6) using high resolution (′′) 11.7 m data from Gemini (see also SOMA II).
From our ATCA observations, we have detected two sources that we name ATCA1 and ATCA2 (see Figure 1). Source ATCA1 is detected at all frequencies (5.5, 9, 17, and 22.8 GHz) and source ATCA2 is only detected at the lower frequencies (5.5 and 9 GHz). Both sources are consistent with the previously reported sources by C. J. Phillips et al. (1998) at 8.6 GHz, sources that are associated with the MIR sources 1 and 2. T. Murphy et al. (2010) presented radio continuum observations at 20 GHz towards this region, detecting only source ATCA1, which is consistent with our results of non-detection for source ATCA2 at higher frequencies.
@@
For ATCA1, T. Murphy et al. (2010) reported a spectral index of 1.2 between 0.843 and 20 GHz, and A. L. Patel et al. (2025) reported a spectral index of 0.5 between 5 and 24 GHz. Our estimated value (0.7) is similar to the one reported by A. L. Patel et al. (2025) at similar frequencies. The estimated spectral index value is consistent with either a typical UC H II region (I. E. Kalcheva et al., 2018) or collimated ionized jets (S. P. Reynolds, 1986; S. J. D. Purser et al., 2016, 2021). Based on our results, we classify this source as an optically thick thermal emitter. However, we cannot rule out the possibility of a radio jet, since masers and molecular outflows have been reported toward this source, and a nearby non-thermal component is present. ATCA2, on the other hand, exhibits non-thermal emission, with an upper limit of , and shows slight extension, as seen in Fig. 1. The nature of this source remains uncertain: given the large distance to this region and the observed morphology, an extragalactic origin cannot be ruled out. Alternatively, this source could represent a jet lobe if ATCA2 was indeed a radio jet. Additional information is required to make a more definitive assessment of the nature of this source.
@@
G35.58-0.03 (hereafter G35.58) is a star-forming region located at a kinematic distance of 10.2 kpc (C. Watson et al., 2003). G35.58 has been resolved into two UC HII regions (from now on called sources 30 and 31) using 2 and 3.6 cm observations, where water and OH masers have been detected towards source 30 (S. Kurtz et al., 1994; J. L. Caswell et al., 1995). Additionally, C.-P. Zhang et al. (2014) shows evidence of an ionized outflow driving a NE-SW molecular outflow centered in source 30.
@@
| 30A | 18:56:22.52 | +02:20:30.5 | 0.7(0.1) | IR | Yesa |
| 30B | 18:56:22.62 | +02:20:34.7 | -1.1 | Yesb | |
| 30C | 18:56:22.57 | +02:20:24.3 | -0.3 | Yesb | |
| 30D | 18:56:22.47 | +02:20:24.8 | -1.2 | Yesb | |
| 30E | 18:56:22.55 | +02:20:21.4 | -0.5(0.1) | Yesb | |
| 30F | 18:56:22.43 | +02:20:20.5 | -0.5 | Yesb | |
| 30G | 18:56:22.39 | +02:20:19.7 | -1.0 | Yesb | |
| 30H | 18:56:22.46 | +02:20:18.9 | -1.9 | Yesb | |
| 30J | 18:56:22.19 | +02:20:21.5 | 0.6 | Yesb | |
| 30K | 18:56:22.19 | +02:20:21.5 | 0.1 | Yesb | |
| 30N | 18:56:22.53 | +02:20:27.6 | -0.5(0.4) | IR, mm | Yesc |
| 30S | 18:56:22.52 | +02:20:27.2 | 0.2(0.1) | IR, mm | Yesc |
| 31 | 18:56:22.62 | +02:20:27.9 | 0.0(0.1) | MIR, mm | No |
@@
G35.58-0.03 30 and 31: Sources 30 and 31 have previously been referred to as UC HII regions (e.g., S. Kurtz et al., 1994). For source 31, based on its morphology in the centimeter and its estimated spectral index, we agree with its nature being an UC HII region. Moreover, source 31 appears to have already a central protostar based on its association with a bright and compact MIR source (SOMA II).
From our observations, source 30 shows a double peak at C-band and a single peak at K-band. At lower frequencies, we marginally resolve it into two components, labeled 30N and 30S (Figure 3), aligned roughly NE–SW, consistent with the larger-scale molecular outflow previously observed in the region. This together with their association with water masers suggests that 30N and 30S could trace an ionized jet. The estimated spectral index of 0.7, derived from the inner scale encompassing source 30, is consistent with typical values for ionized jets associated with young stellar objects (YSOs; e.g., S. P. Reynolds 1986; G. Anglada et al. 1998; K. E. I. Tanaka et al. 2016). In Table 6, we also provide estimates of the spectral indices for 30N and 30S. However, higher-resolution observations will be required to confirm the jet scenario. Finally, the association with multiple star formation tracers further suggests that source 30 may be in a younger evolutionary stage than source 31.
@@ 3.1.4 IRAS 16562-3959
IRAS 16562-3959, also known as G345.49+1.47, is located at a distance of 1.7 kpc and hosts a massive core containing a massive star in an early stage of evolution (A. E. Guzmán et al., 2010). A. E. Guzmán et al. (2011) reported a quadrupolar molecular outflow traced by CO. This morphology is explained by the presence of two collimated bipolar outflows, one lying in the SE–NW direction related to the ionized jet detected by A. E. Guzmán et al. (2010), and another in the N–S direction that could be related to the unresolved millimeter source 13 in A. E. Guzmán et al. (2014).
@@
For our analysis we used A. Guzman’s ATCA observations at 5.5, 9, 17, and 19 GHz from A. E. Guzmán et al. (2016), where they reported nine radio detections associated with this region. The central source has a counterpart at 3 mm and X-ray (source 10 at 3 mm and source 161 in X-ray) (A. E. Guzmán et al., 2014; V. A. Montes et al., 2020), and has also been detected in the MIR by the SOMA II study.
@@
We focus on the regions IRAS 16562-3959 and IRAS 16562-3959 N. The latter was added to the sample in SOMA III due to the detection of a point source in the IR maps, based on the observations analyzed in SOMA II.
@@
IRAS 16562-3959: For the purpose of this study, we only focus on the sources within the SOMA scale (see Figure 1). Detections C, I-E, and I-W all exhibit compact morphologies. As described by A. E. Guzmán et al. (2014), this region shows characteristics consistent with an ionized radio jet, driven by the central free–free emission source (C) and accompanied by two lobes (I-E and I-W) of nonthermal emission. Given the clear jet-like morphology we also measure the fluxes using an intermediate scale, as previously described, that encompass the central source (detection C) and the two inner radio lobes.
@@
Sources 1, 2, and 3 exhibit compact morphologies. A. E. Guzmán et al. (2016) discuss the nature of these detections in Section 4.2.2 of their paper. In summary, Source 1 may be extragalactic, Source 2 is likely an HC H II region associated with a young high-mass star, and the free–free emission from Source 3 originates from a low-mass YSO.
@@
Based on our criteria, and for completeness, we report all detections above a threshold, adopting the same signal-to-noise definition as A. E. Guzmán et al. (2016). As a result, we identify a new source, source 4 (following the convention of A. E. Guzmán et al. 2016), which is detected only at 5 and 9 GHz, shows a slight elongation to the south (Figure 1), and was not previously reported. We estimate an upper limit to its spectral index of .
@@
From previous radio continuum observations toward G305.200.21 there are reports of two detections, which are prominent extended HII regions labeled as G305HII(SE) and G305HII by A. J. Walsh et al. (2007). These regions are located about 30′′ to the southeast of G305A and 15′′ to the southwest of G305B, respectively (A. J. Walsh et al., 2007; C. J. Phillips et al., 1998).
@@
G305A (G305.21+0.21): This source is prominent at MIR and FIR (SOMA II) and has been observed at 8.6 and 18 GHz, but no radio detections have been reported at sensitivity levels of 0.9 mJy and 0.22 mJy, respectively (C. J. Phillips et al., 1998; A. J. Walsh et al., 2007). With our ATCA observations at 5.5, 9, 17 and 22.8 GHz, we report a single detection within the region G305A that we label as G305A ATCA-1 (see Figure 1: G305.21+0.21 (G305A)). This source coincides with the methanol maser reported by A. J. Walsh & M. G. Burton (2006) and shows a slight elongation at the lower frequencies. A. Y. Yang et al. (2021, 2022) reported CO molecular outflows towards this region in the ATLASGAL clump scale.
For G305A we report an estimated spectral index of 0.3, a result that indicates thermal emission. SOMA II suggested that G305A is a much younger and more embedded source than G305B and in a hot core phase, rich in molecular tracers (A. J. Walsh et al., 2007). Considering the spectral index value and star-formation tracers such as reported masers and molecular outflows, we suggest that this source is most likely a radio jet.
G305B (G305.20+0.21): Region G305B is the brightest MIR source and it is associated with methanol masers. However, no radio emission has been previously detected towards this region (A. J. Walsh et al., 2007). In our ATCA observations (see Figure 1) we report 4 sources labeled G305B ATCA-1, G305B ATCA-2, G305B ATCA-3 and G305B ATCA-4 (hereafter sources 1, 2, 3 and 4). In Table 7 we present more information about the multiplicity in G305B such as the R.A. and Decl. positions, the estimated spectral index, association with other wavelengths, and whether this source is a new radio detection or not.
@@
| G305B ATCA-1N | 13:11:10.58 | 62:34:37.8 | 0.4 | IR | Yesa |
| G305B ATCA-1S | 13:11:10.56 | 62:34:38.8 | -1.0 | IR | Yesa |
| G305B ATCA-2 | 13:11:10.04 | 62:34:40.1 | -0.8(0.2) | Yesb | |
| G305B ATCA-3 | 13:11:09.63 | 62:34:38.7 | -0.6(0.2) | Yesb | |
| G305B ATCA-4 | 13:11:10.99 | 62:34:37.8 | -1.7 | Yesb |